The temperature coefficient of resistance for carbon at 0℃ is -0.00048/℃.What is the significance of the minus sign? A carbon resistor has a resistance of 500 Ω at ℃. Determine it's resistance at 50 ℃?


1
Expert's answer
2021-03-03T11:30:08-0500

Explanations & Calculations


  • The temperature dependence of the resistance of a material is given by the equation "\small R_{\theta}=R_0(1+\alpha\Delta\theta)"
  • The temperature coefficient of the resistance determines the behavior of the resistance over the temperature change.
  • Therefore, negativity in that says the resistance of the material is reduced over the increment of temperature.
  • This is the common behavior of the semi-conductor materials like carbon/silicon etc.
  • Using the equation the needed resistance can be found

"\qquad\qquad
\begin{aligned}
\small R_{50}&=\small R_0(1-\alpha\Delta\theta)\\
&=\small500\Omega\big(1-0.00048C^{-1}\times[50C^0-0C^0]\big)\\
&=\small \bold{488\Omega}
\end{aligned}"